OpenStudy (learningallday):

(1 pt) What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (youngstudier):

I'm not sure about this one, I suck at these....

OpenStudy (learningallday):

heres an nother one

OpenStudy (learningallday):

Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5)

OpenStudy (youngstudier):

Remember, the Associative property of Addition is: (a + b) + c = a + (b + c)

OpenStudy (youngstudier):

So what do you think it is?

OpenStudy (youngstudier):

@learningallday

OpenStudy (learningallday):

C. (–8 + 10) + 4. = –8 + (10 + 4)

OpenStudy (learningallday):

@YoungStudier

OpenStudy (youngstudier):

I do believe you are correct.

OpenStudy (learningallday):

what about this one

OpenStudy (learningallday):

What is the value of the expression? –15.6 + (–3.4 + 8) A. –27 B. –26 C. –11 D. –10

OpenStudy (learningallday):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (youngstudier):

-3.4 + 8 = 4.6 -15.6 + 4.6 = -11

OpenStudy (youngstudier):

It would be positive 3 because there is a negative sign outside of the parentheses.

OpenStudy (youngstudier):

So C.

OpenStudy (youngstudier):

Anymore?

OpenStudy (learningallday):

What value of x makes the equation true? –(9 + (–17)) = –9 + x x =

OpenStudy (learningallday):

yup

OpenStudy (learningallday):

@YoungStudier are you here

OpenStudy (youngstudier):

yup, 1 sec

OpenStudy (youngstudier):

Positive 17.

OpenStudy (youngstudier):

Are there anymore?

OpenStudy (learningallday):

thanks i will tell you what i got

OpenStudy (youngstudier):

Hopefully at least 80%.

OpenStudy (learningallday):

80%

OpenStudy (youngstudier):

Hurray! Which one did you get wrong? The 1st one?

OpenStudy (learningallday):

yes

OpenStudy (youngstudier):

Eh, well, you still passed is all that matters. :D

OpenStudy (learningallday):

what grade are you in

OpenStudy (youngstudier):

7th grade, doing 8th grade math.

OpenStudy (learningallday):

cool

OpenStudy (youngstudier):

Yup

OpenStudy (learningallday):

me to

OpenStudy (learningallday):

im in k12

OpenStudy (youngstudier):

You're doing 8th grade math, or you're in 7th grade?

OpenStudy (learningallday):

i dont know